# LZ Penalty: An information-theoretic repetition penalty for autoregressive language models

### v2

@@ Abstract
We introduce the Lemepel-Ziv (LZ) penalty, a penalty specialized for reducing degenerate repetitions in autoregressive language models without loss of capability. The penalty is based on the codelengths in the LZ77 universal lossless compression algorithm. Through the lens of the prediction-compression duality, decoding with the LZ penalty has the interpretation of sampling from the residual distribution after removing the information that is highly compressible. We demonstrate the LZ penalty enables open-source reasoning models to operate with greedy decoding without loss of capability and without instances of degenerate repetition. Both the industry-standard frequency penalty and repetition penalty are ineffective, incurring degenerate repetition rates of up to $4\%$ or more.
@@ 1 Introduction
In recent months, there has been an advent in reasoning models. Reasoning models are a class of large, autoregressive foundation models that achieve impressive capability gains in certain domains by scaling chain-of-thought reasoning sequences at inference time. While reasoning models are a promising approach for scaling inference-time compute, open-source reasoning models currently suffer from some friction points that make their use problematic for downstream application developers due to a lack of determinism around the reasoning traces. Open-source reasoning models do not run well at low temperatures because the sampling distribution mode collapses into degenerate repetitions.
Enabling deterministic algorithms for generation is useful for debugging and may, in fact, be explicit desiderata for some deployments.
There are two industry-standard penalties aimed at reducing repetition. First, the repetition penalty [7] applies a fixed logit penalty that encourages the model to use new tokens. The frequency penalty, is more subtle, and applies a logit penalty proportional to the token count in context. Neither penalty consistently stops degenerate repetitions without degrading the sample quality. First, the repetition penalty does not actually succeed in preventing degenerate repetitions because it applies a naive, binary modal penalty which does not take into account the number of times a token has appeared. Furthermore, if the repetition penalty is set too high in an effort to minimize this mode collapse, the sampler becomes unable to use fundamental, necessary, but frequent tokens, such as spaces or periods, resulting in poor completions. Thus, merely increasing the repetition penalty is not a viable solution either.
@@
The fundamental improvement in the LZ penalty relative to the repetition or frequency penalty is that the LZ penalty, borrowing from the sliding window matching techniques pioneered in the LZ77 [8] and LZSS [18] lossless compression algorithms, depends on the repetition of $n$-grams over a long but fixed-length sliding window. By penalizing as a function of length-normalized $n$-gram statistics as opposed to single token statistics, the penalty can be significantly more surgical in how it modulates the sampling distribution.
@@
Summary of Contributions
We propose the LZ penalty, a novel sampling penalty that demonstrates excellent mitigation of degenerate repetition with greedy sampling in frontier open-source models without loss of capability on difficult reasoning benchmarks such as AIME and GPQA. For practitioners, we provide a batteries-included reference implementation [^2] https://github.com/tginart/sglang, https://github.com/tginart/lz-penalty in SGLang. While not fully optimized, the reference implementation incurs a slowdown of less than $0.03\%$ in median throughput and no measurable slowdown in median latency.
Limitations
There are two main limitations to the LZ penalty. First, it is somewhat more complex than the simpler frequency penalty. Second, the LZ penalty addresses determinism at the algorithmic-level, but algorithmic determinism is not always necessary or sufficient in practice. At the application-level, persistent caching or fixed seeding could be used to ensure the same input context consistently produces the same completion. Furthermore, hardware-level non-determinism is possible on certain accelerators, potentially defeating determinism at the algorithmic-level.
@@ 2.1 Language Modeling and Sampling
Language model traces its origins back to Shannon’s testament [17], where he trained a casual language model by computing the $n$-gram frequencies over an English text corpus. Modern language models are predominantly based on the transformer architecture [20]. Completions from transformer language models are generated by autoregressive sampling of the next-token distribution. The development of transformer language models has been accompanied by significant advancements in sampling techniques that govern text generation. Early explorations of language modeling employed top-$k$ sampling [6] to constrain the output distribution to the $k$ most probable tokens, a technique later refined in [22], which paired it with unlikelihood training to mitigate repetition. Concurrently, temperature sampling [1] emerged as a method to control randomness. Later, nucleus sampling was introduced as a proposed improvement over top-$k$ sampling [4].
These sampling strategies evolved alongside efforts to address text degeneration and repetition. GPT-2’s implementation [15] implicitly utilized frequency penalties to enhance output fluency, and, concurrently the repetition penalty [7] was devised to prevent repetitions and encourage diversity in completions. Later, LaMDA [19] applied repetition penalties to improve dialog coherence, reflecting a growing emphasis on balancing creativity and quality in LLM outputs. Together, these contributions and others eventually led to an industry-standard sampler which supports a temperature, a top-$k$, a top-$p$ and a frequency or repetition penalty, all of which can be used in tandem to transform the raw next token logits into a final distribution for sampling.
@@ 2.2 Data Compression
Data compression algorithms go back to the turn of the 20th century. In Shannon’s testament [17], he describes the first provably optimal compressor. Later, many entropy-optimal compressors achieved practical computational complexity assuming known data distributions. Later still, in 1977 and 1978, the first universal compressors were launched, LZ77 and LZ78, that could, asymptotically, achieve the entropy-rate of any stationary, ergodic data source [8, 9, 24, 11]. Since then, a whole family of LZ-style compressors has emerged [3, 10, 14, 12, 25, 18, 21].
@@
Single-Token Data Compressor. A single-token data compressor, $\mathcal{C}:V\rightarrow\{0,1\}^{*}$ maps literal single-tokens to binary codes. We will assume single-token data compressors are complete prefix codes [2].
@@
In this work, we will only focus on the LZ77 family, which we refer to herein as the LZ sliding window compression algorithm, which uses string matching from a buffer over a lookback window. This contrasts the LZ78 family, which favors tree-style dictionaries. Sliding windows are more convenient for GPUs (for example, by using PyTorch’s [13] unfold operation) whereas tree-based dictionary methods are more inherently sequential.
@@
Concretely, LZSS [18] modifies LZ77 by using a 1-bit flag to indicate whether the next chunk of data is a literal or a length-distance pair and uses literals if a length–distance pair is below a given minimum length. Since we do not actually need to encode or decode the token sequence, the details of the encoding subroutine are not particularly important for our purposes. Instead, we should focus on how many bits are required for the encodings — the codelengths of the resulting codes. We take LZSS as our reference compressor herein, and use the LZSS encoding scheme in the LZ penalty. When we refer to a generic LZ sliding window algorithm, we will mean the LZSS variant.
@@
$\mathcal{C}_{LZ}(\mathbf{b}|\mathbf{w})=\begin{cases}\mathcal{C}(d,l)&\text{if}l\geq 1\textbf{ and }l=|\mathbf{b|}\\ \mathcal{C}(d,l)\mathcal{C}_{LZ}(\mathbf{b}_{>l}|\mathbf{w})&\text{if }l\geq 1\textbf{ and }l<|\mathbf{b|}\\ \mathcal{C}(\mathbf{b}_{1})\mathcal{C}_{LZ}(\mathbf{b}_{>1}|\mathbf{w})&\text{if }l=0\\ \end{cases}$
@@
[18] LZSS can encode a match of length $L$ occurring $D$ tokens in the past using $|\mathcal{C}_{LZ}(L,D)|=\log L+\log D+1$ bits.
@@
[18] LZSS requires $|\mathcal{C}_{LZ}(a)|=\log V+1$ bits to encode token literals $a\in V$.
@@ 2.3 The Prediction-Compression Duality
Let $\mathbf{DC}$ be a (prefix-free) single-token compressor. Define the logits of a dual language model as:
[…]
@@
[…]
@@
Then, the Arithmetic coding construction ([2, 23]) produces a sequential prefix-free compressor $\mathcal{C}_{\mathbf{LM}}$ satisfying, for every $x\in V^{n}$,
@@
[…]
@@
Given a language model, we also have a data compressor that compresses as well as the language model predicts, and given a data compressor, we have a language model that predicts as well as that data compressor can compress. The Arithmetic code (and other codes such as the Huffman code [5, 2]), employ the prediction-compression duality to assign codelengths based on log-probabilities.
The situation is more complex for constructing casual language models from online data compressors such as LZ sliding window algorithms. This is because casual language models must be able to generate a valid next-token pmf at every step whereas data compressors often buffer tokens together into a single code. Practically, this means data compressors do not necessarily produce a codelength for every next-token. We address this issue by simulating a full buffer and lookback window at each next-token. Similar ideas have been explored in [16].
@@ 3 LZ Penalty
$\lambda(a)=\begin{cases}1&\text{if }a\in\mathbf{w}(x)\text{ and }l=0\\ 0&\text{if }a\notin\mathbf{w}(x)\text{ and }l=0\\ \end{cases}$
@@
$\mathcal{C}_{\mathrm{I}}=\mathcal{C}(a|\mathbf{w})=\begin{cases}\mathcal{C}(1,\delta)&\text{if }\lambda=1\\ \mathcal{C}(a)&\text{if }\lambda=0\\ \end{cases}$ (7)
@@
$|\mathcal{C}_{\mathrm{I}}|=\begin{cases}\log\delta&\text{if }\lambda=1\\ \log V&\text{if }\lambda=0\\ \end{cases}$ (8)
@@
$\mathcal{C}_{\mathrm{II}}=\mathcal{C}(a\mathbf{b}_{\leq l}|\mathbf{w})=\begin{cases}\mathcal{C}(a|\mathbf{w})\mathcal{C}(\mathbf{b}_{\leq l}|\mathbf{w})=\mathcal{C}(a|\mathbf{w})\mathcal{C}(l,d)&\text{if }\lambda\leq 1\\ \mathcal{C}(a\mathbf{b}_{\leq l}|\mathbf{w})=\mathcal{C}(\lambda,\delta)&\text{if }\lambda=l+1\\ \end{cases}$ (10)
@@
$\mathcal{C}_{\mathrm{II}}=\begin{cases}\mathcal{C}(a)\mathcal{C}(l,d)&\text{if}\lambda=0\\ \mathcal{C}(1,\delta)\mathcal{C}(l,d)&\text{if }\lambda=1\\ \mathcal{C}(\lambda,\delta)&\text{if }\lambda=l+1\\ \end{cases}$ (11)
@@
$|\mathcal{C}_{\mathrm{II}}|=\begin{cases}|\mathcal{C}(a)\mathcal{C}(l,d)|=|\mathcal{C}(a)|+|\mathcal{C}(l,d)|=\log V+\log(ld)+1&\text{if }\lambda=0\\ |\mathcal{C}(1,\delta)|+|\mathcal{C}(l,d)|=\log(\delta)+\log(ld)+1&\text{if }\lambda=1\\ |\mathcal{C}(\lambda,\delta)|=\log(\lambda\delta)&\text{if }\lambda=l+1\\ \end{cases}$ (12)
@@
$|\mathcal{C}_{\mathrm{II}}|=\begin{cases}\log V&\text{if }\lambda=0\\ \log(\delta)&\text{if }\lambda=1\\ \log(1-\frac{d-l+1}{ld})-1&\text{if }\lambda=l+1\\ \end{cases}$ (13)
@@
$\Delta|\mathcal{C}_{LZ}|=\begin{cases}\log V&\text{if }\lambda=0\\ \log(\delta)&\text{if }\lambda=1\\ \log(1-\frac{d-l+1}{ld})-1&\text{if }\lambda=l+1\\ \end{cases}$ (14)
Assuming $V>|\mathbf{w}|$, then this provides a dynamic range of $[\log(2/|\mathbf{b}|)-1,\log V]$. For a vocabulary of size $128k$, a lookback window of size $512$, and a buffer of size $32$, using binary logarithms, this yields on adjustment range from $-5$ to $+17$, with a $-5$ adjustment going to a token that would complete an immediate repetition of length $32$ and a $+17$ going to a token that does not appear in the previous $512$ tokens.
@@ 4.2 Latency and Throughput Benchmark
While our SGLang reference implementation is not fully optimized, it it is vectorized and batched. We run SGLang’s built-in benchmark script on an 8×8$\times$8 ×H100 node and compare the effect of adding non-zero LZ penalty. While the LZ penalty adds an ultimately negligible amount of computation, it still is significantly more than, say, the repetition penalty, so this it is worthwhile to confirm that we can maintain inference performance.
We see that for larger models, the LZ penalty’s overhead is increasingly negligble. Even for models as small as $1.5$B, the penalty overhead is a tolerable $0.55$% throughput slowdown. For latency, the overhead is more trivial, and is not even measurable at the $32$B size.
@@ Appendix A Appendix
A.1 Supplementary Results
@@
As mentioned in the main text, we include sweeps for the penalty strengths for both the repetition and the frequency penalty. For the repetition penalty, we sweep $\{1,1.1,1.2,1.25,1.3,1.5\}$ with $1$ being the default value for a no-penalty baseline and $1.2$ being a standard value suggest by [7]. For the frequency penalty, we sweep $\{0,0.1,0.2,0.3,0.4,0.6,0.8,0.9,1.0\}$ with 00 being the default value for a no-penalty baseline.
Figure 4: A heatmap comparing accuracy and repetition percentage for the LZ penalty and frequnecy penalty across temperatures and penalty strengths. Each row reports a temperature sweep over a fixed frequency penalty strength, except for the first row, which does not use the standard frequency penalty and instead uses the proposed LZ penalty. Hence the LZ label on the first row.

### v3

@@ Abstract
We introduce the Lempel-Ziv (LZ) penalty, a penalty specialized for reducing degenerate repetitions in autoregressive language models without loss of capability. The penalty is based on the codelengths in the LZ77 universal lossless compression algorithm. Through the lens of the prediction-compression duality, decoding with the LZ penalty has the interpretation of sampling from the residual distribution after removing the information that is highly compressible. We demonstrate the LZ penalty enables open-source reasoning models to operate with greedy decoding without loss of capability and without instances of degenerate repetition. Both the industry-standard frequency penalty and repetition penalty are ineffective, incurring degenerate repetition rates of up to $4\%$ or more.
@@ 1 Introduction
In recent months, there has been an advent in reasoning models. Reasoning models are a class of large, autoregressive foundation models that achieve impressive capability gains in certain domains by scaling chain-of-thought reasoning sequences at inference time. While reasoning models are a promising approach for scaling inference-time compute, open-source reasoning models currently suffer from some friction points that make their use problematic for downstream application developers due to a lack of determinism around the reasoning traces. Reasoning models do not run well at low temperatures because the sampling distribution mode collapses into degenerate repetitions.
Enabling deterministic algorithms for generation is useful for debugging and may, in fact, be explicit desiderata for some deployments. Furthermore, even at higher temperatures, even frontier models can still fall into degenerate repetitions in real-world deployments, such as within Cursor [^1] See Appendix. www.cursor.com..
There are two industry-standard penalties aimed at reducing repetition. First, the repetition penalty (Keskar et al., 2019) applies a fixed logit penalty that encourages the model to use new tokens. The frequency penalty, is more subtle, and applies a logit penalty proportional to the token count in context. Neither penalty consistently stops degenerate repetitions without degrading the sample quality. First, the repetition penalty does not actually succeed in preventing degenerate repetitions because it applies a naive, binary modal penalty which does not take into account the number of times a token has appeared. Furthermore, if the repetition penalty is set too high in an effort to minimize this mode collapse, the sampler becomes unable to use fundamental, necessary, but frequent tokens, such as spaces or periods, resulting in poor completions. Thus, merely increasing the repetition penalty is not a viable solution either.
@@
The fundamental improvement in the LZ penalty relative to the repetition or frequency penalty is that the LZ penalty, borrowing from the sliding window matching techniques pioneered in the LZ77 (Lempel & Ziv, 1977) and LZSS (Storer & Szymanski, 1982) lossless compression algorithms, depends on the repetition of $n$-grams over a long but fixed-length sliding window. By penalizing as a function of length-normalized $n$-gram statistics as opposed to single token statistics, the penalty can be significantly more surgical in how it modulates the sampling distribution.
@@ 2.1 Language Modeling and Sampling
Language model traces its origins back to Shannon’s testament (Shannon, 1948), where he trained a casual language model by computing the $n$-gram frequencies over an English text corpus. Modern language models are predominantly based on the transformer architecture (Vaswani et al., 2017). Completions from transformer language models are generated by autoregressive sampling of the next-token distribution. The development of transformer language models has been accompanied by significant advancements in sampling techniques that govern text generation. Early explorations of language modeling employed top-$k$ sampling (Jozefowicz et al., 2016) to constrain the output distribution to the $k$ most probable tokens, a technique later refined in (Welleck et al., 2019), which paired it with unlikelihood training to mitigate repetition. Concurrently, temperature sampling (Bowman et al., 2016) emerged as a method to control randomness. Later, nucleus sampling was introduced as a proposed improvement over top-$k$ sampling (Holtzman et al., 2020).
These sampling strategies evolved alongside efforts to address text degeneration and repetition. GPT-2’s implementation (Radford et al., 2019) implicitly utilized frequency penalties to enhance output fluency, and, concurrently the repetition penalty (Keskar et al., 2019) was devised to prevent repetitions and encourage diversity in completions. Later, LaMDA (Thoppilan et al., 2022) applied repetition penalties to improve dialog coherence, reflecting a growing emphasis on balancing creativity and quality in LLM outputs. Together, these contributions and others eventually led to an industry-standard sampler which supports a temperature, a top-$k$, a top-$p$ and a frequency or repetition penalty, all of which can be used in tandem to transform the raw next token logits into a final distribution for sampling.
@@ 2.2 Data Compression
Data compression algorithms go back to the turn of the 20th century. In Shannon’s testament (Shannon, 1948), he describes the first provably optimal compressor. Later, many entropy-optimal compressors achieved practical computational complexity assuming known data distributions. Later still, in 1977 and 1978, the first universal compressors were launched, LZ77 and LZ78, that could, asymptotically, achieve the entropy-rate of any stationary, ergodic data source (Lempel & Ziv, 1977; 1978; Wyner & Ziv, 1994; Morita & Kobayashi, 1993). Since then, a whole family of LZ-style compressors has emerged (Fiala & Greene, 1989; Miller & Wegman, 1985; Pavlov, 2007; Oberhumer, 1997; Yoshizaki, 1988; Storer & Szymanski, 1982; Welch, 1984).
@@
Single-Token Data Compressor. A single-token data compressor, $\mathcal{C}:V\rightarrow\{0,1\}^{*}$ maps literal single-tokens to binary codes. We will assume single-token data compressors are complete prefix codes (Cover & Thomas, 2006).
@@
In this work, we will only focus on the LZ77 family, which we refer to herein as the LZ sliding window compression algorithm, which uses string matching from a buffer over a lookback window. This contrasts the LZ78 family, which favors tree-style dictionaries. Sliding windows are more convenient for GPUs (for example, by using PyTorch’s (Paszke et al., 2019) unfold operation) whereas tree-based dictionary methods are more inherently sequential.
@@
Concretely, LZSS (Storer & Szymanski, 1982) modifies LZ77 by using a 1-bit flag to indicate whether the next chunk of data is a literal or a length-distance pair and uses literals if a length–distance pair is below a given minimum length. Since we do not actually need to encode or decode the token sequence, the details of the encoding subroutine are not particularly important for our purposes. Instead, we should focus on how many bits are required for the encodings — the codelengths of the resulting codes. We take LZSS as our reference compressor herein, and use the LZSS encoding scheme in the LZ penalty. When we refer to a generic LZ sliding window algorithm, we will mean the LZSS variant.
@@
[…]
@@
(Storer & Szymanski, 1982) LZSS can encode a match of length $L$ occurring $D$ tokens in the past using $|\mathcal{C}_{LZ}(L,D)|=\log L+\log D+1$ bits.
@@
(Storer & Szymanski, 1982) LZSS requires $|\mathcal{C}_{LZ}(a)|=\log V+1$ bits to encode token literals $a\in V$.
@@ 2.3 The Prediction-Compression Duality
Let $\mathbf{DC}$ be a single-token compressor. Define the logits of a dual language model as:
$\ell_{\mathbf{DC}}\!\bigl(x_{i}\mid x_{<i}\bigr)\;:=\;\bigl|\mathcal{C}_{\mathbf{DC}}\!\bigl(x_{i}\mid x_{<i}\bigr)\bigr|\qquad\text{(bits)}$
@@
$\bigl|\bar{\mathcal{C}}_{\mathbf{DC}}\bigr|(x)\;=\;\frac{1}{n}\sum_{i=1}^{n}\ell_{\mathbf{DC}}\!\bigl(x_{i}\mid x_{<i}\bigr)\;=\;-\frac{1}{n}\sum_{i=1}^{n}\log p_{\mathbf{DC}}\!\bigl(x_{i}\mid x_{<i}\bigr)\;=\;H_{\mathbf{DC}}(x)\;\;\text{bits/token}.$
@@
Then, the Arithmetic coding construction ((Cover & Thomas, 2006; Witten et al., 1987)) produces a sequential prefix-free compressor $\mathcal{C}_{\mathbf{LM}}$ satisfying, for every $x\in V^{n}$,
@@
$H_{\mathbf{LM}}(x)\;+\;O\!\bigl(1/n\bigr)\;\text{bits/token}.$
@@
Given a language model, we also have a data compressor that compresses as well as the language model predicts, and given a data compressor, we have a language model that predicts as well as that data compressor can compress. The Arithmetic code (and other codes such as the Huffman code (Huffman, 1952; Cover & Thomas, 2006)), employ the prediction-compression duality to assign codelengths based on log-probabilities.
The situation is more complex for constructing casual language models from online data compressors such as LZ sliding window algorithms. This is because casual language models must be able to generate a valid next-token pmf at every step whereas data compressors often buffer tokens together into a single code. Practically, this means data compressors do not necessarily produce a codelength for every next-token. We address this issue by simulating a full buffer and lookback window at each next-token. Similar ideas have been explored in (Ryabko, 2007).
@@ 3 LZ Penalty
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
Assuming $V>|\mathbf{w}|$, then this provides a dynamic range of $\log(2/|\mathbf{b}|)-1,\log V]$. For a vocabulary of size $128k$, a lookback window of size $512$, and a buffer of size $32$, using binary logarithms, this yields on adjustment range from 5-5 to +17+17, with a 5-5 adjustment going to a token that would complete an immediate repetition of length $32$ and a +17+17 going to a token that does not appear in the previous $512$ tokens.
@@ 4.2 Latency and Throughput Benchmark
While our SGLang reference implementation is not fully optimized, it it is vectorized and batched. We run SGLang’s built-in benchmark script on an $8\timesH$100 node and compare the effect of adding non-zero LZ penalty. While the LZ penalty adds an ultimately negligible amount of computation, it still is significantly more than, say, the repetition penalty, so this it is worthwhile to confirm that we can maintain inference performance.
We see that for larger models, the LZ penalty’s overhead is increasingly negligble. Even for models as small as 1.51.5B, the penalty overhead is a tolerable 0.550.55% throughput slowdown. For latency, the overhead is more trivial, and is not even measurable at the 3232B size.
@@ Appendix A Appendix
A.1 Evidence of Degenerate Repetitions
Although uncommon, degenerate repetitions do still occur in category-leading AI services powered by frontier AI models. Here we provide documentation of such an incident in Cursor888www.cursor.com using Gemini-2.5-Pro [^9] https://blog.google/technology/google-deepmind/gemini-model-thinking-updates-march-2025/#gemini-2-5-thinking.
Figure 3: A screenshot of a degenerate repetition in an Cursor agent mode request for Gemini-2.5-Pro.
A.2 Supplementary Results
As mentioned in the main text, we include sweeps for the penalty strengths for both the repetition and the frequency penalty. For the repetition penalty, we sweep $\{1,1.1,1.2,1.25,1.3,1.5\}$ with $1$ being the default value for a no-penalty baseline and $1.2$ being a standard value suggest by Keskar et al. (2019). For the frequency penalty, we sweep $\{0,0.1,0.2,0.3,0.4,0.6,0.8,0.9,1.0\}$ with 0 being the default value for a no-penalty baseline.
Figure 4: A heatmap comparing accuracy and repetition percentage for the LZ penalty and penalty across temperatures and frequency penalty strengths. Each row reports a temperature sweep over a fixed repetition penalty strength, except for the first row, which does not use the standard repetition penalty and instead uses the proposed LZ penalty. Hence the LZ label on the first row.